# Digital Twin Calibration for Biological System-of-Systems: Cell Culture Manufacturing Process

### v1

@@ 6 Empirical Study
    To assess the performance of the proposed calibration approach, we employ a simpler version of the Chinese Hamster Ovary (CHO) cell culture model as described by Ghorbaniaghdam et al. (2014) as our SMRN model. We use this model with a state shift model to simulate cell culture dynamics and compare our gradient-based policy against a random policy for a DoE problem. We detail the model and its validation in Section 6.1, and present the comparative results in Section 6.2.
@@ 6.1 Cell Culture Model and Validation
    CHO cells have become the most commonly used mammalian hosts for industrial production of monoclonal antibodies (mAbs) and recombinant protein therapeutics. This is due to their capacity to correctly fold, assemble, and modify human proteins, ensuring biological activity and stability. The robustness of CHO cells, coupled with their genetic and metabolic malleability, makes them ideal candidates for genetic engineering and optimization for specific production processes. The cell growth model is described by a function characterizing different phases of cell culture
@@
    where $v_{growth0}$ represents the growth rate that depends on glucose concentration in the growth phase (phase 0) and where $\beta_{0}$ and $\beta_{1}$ are phase transition rates, reflecting the dynamic interplay between growth and production phases. A simplified SMRN focuses on three key metabolisms: glucose consumption, lactate production, and monoclonal antibody synthesis with chemical equations
    $GLC$ $\rightarrow 2LAC,\quad 0.02GLC\rightarrow X,\quad 0.01GLC\rightarrow mAb$
    where GLC, LAC and mAb represent glucose, lactate and monoclonal antibodies. The state of the system at any time $t$ is denoted by $\boldsymbol{s}_{t}=\left[[GLC]_{t},[LAC]_{t},[mAb]_{t}\right]$ representing the concentrations of glucose, lactate, and mAb, respectively. The flux rates $\boldsymbol{v}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})=(-v_{HK},2v_{HK},v_{mAb})$, where $-v_{HK}$ denotes the rate of glucose consumption, $2v_{HK}$ represents the stoichiometric production of lactate , and $v_{mAb}$ quantifies the mAb production rate. The kinetic equations of two phases $i=0,1$ are described as
@@
    The metabolic reactions are formulated as a SDE system (14) to account for the inherent randomness and fluctuations in biochemical reactions. The general form of this SDE is given by
    $\mathrm{d}\boldsymbol{s}_{t}(\boldsymbol{\theta})=\boldsymbol{b}(\boldsymbol{s}_{t}(\boldsymbol{\theta}))\text{d}t+\mathbf{Q}(\boldsymbol{s}_{t}(\boldsymbol{\theta}))\text{d}\boldsymbol{B}_{t},$
    where the mean function $\boldsymbol{b}=\mathbf{N}\sum_{i=0}^{1}X_{i,t+\text{d}t}\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$ is modeled by the stoichiometric matrix
    $\mathbf{N}=\begin{bmatrix}-1&-0.02&-0.01\\ 2&0&0\\ 0&0&1\end{bmatrix}$
    and reaction rates $\boldsymbol{v}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. The system variability is modeled by $\mathbf{Q}=\left\{\mathbf{N}\text{diag}\{\sum_{i=0}^{1}X_{i,t+\text{d}t}\boldsymbol{\nu}^{r}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})\}\mathbf{N}^{\top}\right\}^{\frac{1}{2}}$.
    We fit our model by minimizing $\text{MSE}=\frac{1}{n}\sum_{j=1}^{n}{[Y^{s}_{j}-Y^{e}_{j}]^{2}}$ between experiment data $Y^{e}$ from Ghorbaniaghdam et al. (2014) and simulation model prediction $Y^{s}$. The results of our model validation are depicted in Figure 3. The figure presents experimental data as red dots and the simulation results as blue lines, which illustrate both the mean and the 95% confidence interval (CI) of the model predictions. The CI is calculated by the formula $[\bar{x}-z{\sigma}/{\sqrt{n}},\bar{x}+z{\sigma}/{\sqrt{n}}]$, where $\bar{x}$ is the sample mean, $z$ is the z-score corresponding to the 95% confidence level, which is approximately 1.96, $\sigma$ is the standard deviation of the sample, $n$ is the number of observations in the sample. This result shows the model’s effectiveness in capturing the dynamics of the CHO cell culture system.
    Figure 3: The cell culture mechanistic model was validated by using the experimental data from the literature (Ghorbaniaghdam et al., 2014). Red dots represent the experimental data from the literature and the simulation predictions by our model is demonstrated in blue lines with mean and 95% CI computed from $n=20$ replications.
    In the empirical study assessing the performance of the proposed calibration approach, we use the fitted parameters as the true model parameters of the physical system, i.e.,
    $\boldsymbol{\alpha}_{0}$ $=[v_{maxHK0},v_{maxmAb0},K_{mGLC0},v_{maxg0}]=[4.15\times 10^{-5},3.5\times 10^{-8},4.2,0.126]$ (20)
    $\boldsymbol{\alpha}_{1}$ $=[v_{maxHK1},v_{maxmAb1},K_{mGLC1}]=[1.95\times 10^{-5},7.5\times 10^{-8},3.5]$ (21)
    $\boldsymbol{\beta}$ $=[\beta_{0},\beta_{1}]=[0.3,0.1].$ (22)
    and assume the digital twin has the same model structure with unknown parameters.
@@ 6.2 Digital Twin Calibration
    With the true parameters defined in (20) to (22), our objective is to calibrate the digital twin model by strategically designing CHO cell culture experiments. The aim is to minimize the model prediction error by choosing the initial glucose concentration as design decision variable. In each experiment, we adjust the initial glucose concentration and then collect a new batch of data to sequentially refine our model.
    For the calibration experiments setting, we assume time course data of the states is collected every 12h until 72h, so each batch includes 7-time sequence data. We focus on calibrating parameters $\boldsymbol{\alpha}_{0},\boldsymbol{\alpha}_{1}$. The design space for the initial concentrations of glucose is set within the range $[18,38]$ mM, and the initial concentration of lactate and mAb are $[LAC]_{0}=0.8$ mM, $[mAb]_{0}=0$ mM, the initial cell densities in Phase 0 and 1 are set to be $X_{0,0}=180\times 10^{6}$ cells/L, $X_{1,0}=10\times 10^{6}$ cells/L, respectively. The initial parameters for the digital twin are estimated based on the first batch of data. Since the model parameters are in different scales, we set a learning rate $\lambda=[1\times 10^{-5},1\times 10^{-7},1,0.1,1\times 10^{-5},1\times 10^{-7},0.1]$ for updating the calibrating parameters $(\boldsymbol{\alpha}_{0},\boldsymbol{\alpha}_{1})$ and $\gamma_{k}=5$ for updating the design policy by applying Eq. (19). The fixed set of test data $\mathcal{D}_{test}$ is composed of initial states $\boldsymbol{s}_{0}^{1}=[28,0.8,0]$ and $\boldsymbol{s}_{0}^{2}=[38,0.8,0]$.
    We conducted 15 macro-replications, each consisting of $K=200$ iterations under varying initial conditions, with initial glucose levels randomly chosen by applying the uniform distribution over the interval $[18,38]$. We compare the proposed gradient-based calibration approach with a random design approach, where the design is uniformly sampled from $[8,38]$. To ensure consistency across the two approaches, we used the same random seeds for these two approaches. The mean prediction performance and the 95% confidence interval, calculated across macro-replications, represented by the Mean Relative Errors $\text{MRE}=\frac{1}{7}\sum_{j=1}^{7}{|\widehat{\theta}_{(j)}-\theta_{(j)}^{c}|}/{|\theta_{(j)}^{c}|}$ with script $(j)$ representing the $j$-th component, are presented in Figure 5 for mAb prediction and in Figure 5 for parameter estimation, respectively. For both figures, blue lines represent the performance of our approach, while yellow lines represent the random design performance.
    Figure 4: Yield Prediction Error
@@
    The results demonstrate that the proposed Bio-SoS model calibration approach is significantly more sample-efficient than a random design. After 125 iterations, MRE for mAb production predictions using our calibration approach is reduced to approximately 60%, in contrast to roughly 200% for predictions made through random design. Furthermore, the MRE for parameters using our approach decreases to around 30% after 125 iterations, whereas the MRE for the random policy stays as high as 400%.
@@ 7 CONCLUSION
    This study develops a robust digital twin calibration approach for a Bio-SoS mechanistic model in the context of cell culture processes. By strategically guiding more informative data collection through the proposed gradient-based calibration approach, we significantly enhance the model fidelity and predictive accuracy of the digital twin. The empirical validation using the CHO cell culture model underscores the superiority of our approach over the traditional random design, showcasing its potential applicability across various biological systems. Overall, this study not only contributes to the theoretical advancements in bioprocess digital twin development but also holds promise for practical implementations that could improve the efficiency and effectiveness of biomanufacturing in the pharmaceutical industry.

### v2

@@ 6 Empirical Study
    For validation, we use a simplified version of the Chinese Hamster Ovary (CHO) cell culture model (Ghorbaniaghdam et al., 2014) as our SMRN model. We combine it with a state shift model to simulate cell culture dynamics. Our gradient-based policy is compared against a random policy and a GP-based calibration approach. Details of experiment setup are provided in Section 6.1, with results in Section 6.2.
@@ 6.1 Cell Culture Model and Validation
    CHO cells have become the most commonly used mammalian hosts for the industrial production of monoclonal antibodies (mAbs) and recombinant protein therapeutics. The cell growth model is described by a function characterizing different phases of cell culture
@@
    where $v_{growth0}$ represents the growth rate that depends on glucose concentration in the growth phase (phase 0) and where $\beta_{0}$ and $\beta_{1}$ are phase transition rates, reflecting the dynamic interplay between growth and production phases. A simplified SMRN focuses on three key metabolisms: glucose consumption, lactate production, and mAb synthesis with chemical equations $(1)~{}\text{GLC}\rightarrow 2\text{LAC}$, $(2)~{}0.02\text{GLC}\rightarrow\text{X}$, and $(3)~{}0.01\text{GLC}\rightarrow\text{mAb}$, where GLC and LAC represent glucose and lactate. The state of the system at any time $t$ is denoted by $\boldsymbol{s}_{t}=\left[[GLC]_{t},[LAC]_{t},[mAb]_{t}\right]$ representing the concentrations of glucose, lactate, and mAb, respectively. The flux rates $\boldsymbol{v}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})=(-v_{HK},2v_{HK},v_{mAb})$, where
@@
    The metabolic reactions are formulated as a SDE system (12) to account for the inherent randomness and fluctuations in biochemical reactions. The general form of this SDE is given by $\mathrm{d}\boldsymbol{s}_{t}(\boldsymbol{\theta})=\boldsymbol{b}(\boldsymbol{s}_{t}(\boldsymbol{\theta}))\text{d}t+\mathbf{Q}(\boldsymbol{s}_{t}(\boldsymbol{\theta}))\text{d}\boldsymbol{B}_{t},$ where the mean function $\boldsymbol{b}=\mathbf{N}\sum_{i=0}^{1}X_{i,t+\text{d}t}\boldsymbol{\nu}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$ is modeled by the stoichiometric matrix $\mathbf{N}=[-1,-0.02,-0.01;~{}2,0,0;~{}0,0,1]$ and reaction rates $\boldsymbol{v}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})$. The system variability is modeled by $\mathbf{Q}=\left\{\mathbf{N}\text{diag}\{\sum_{i=0}^{1}X_{i,t+\text{d}t}\boldsymbol{\nu}^{r}(\boldsymbol{s}_{t};\boldsymbol{\alpha}_{i})\}\mathbf{N}^{\top}\right\}^{\frac{1}{2}}$.
    We fit our model by minimizing $\text{MSE}=\frac{1}{n}\sum_{j=1}^{n}{[Y^{s}_{j}-Y^{e}_{j}]^{2}}$ between experiment data $Y^{e}$ from Ghorbaniaghdam et al. (2014) and simulation $Y^{s}$ from our model. The results of our model validation are depicted in Figure 3, where the measurements are from Ghorbaniaghdam et al. (2014). The confidence interval (CI) is calculated by the formula $[\bar{x}-z{\sigma}/{\sqrt{n}},\bar{x}+z{\sigma}/{\sqrt{n}}]$, where $\bar{x}$ is the sample mean, $z$ is the z-score corresponding to the 95% confidence level, which is approximately 1.96, $\sigma$ is the standard deviation of the sample, $n$ is the number of observations in the sample. This result shows the model’s effectiveness in capturing the dynamics of the CHO cell culture system.
    Figure 3: The cell culture mechanistic model was validated by using the experimental data. Red dots represent the experimental data from the literature, while the blue lines indicate the simulation predictions from our model, with the mean and 95% CI calculated from 20 replications.
    In the following calibration experiments, we use the fitted parameter as the true model parameter of the physical system and assume the digital twin has the same model structure with unknown parameters. The true parameters are (1) $\boldsymbol{\alpha}_{0}=[v_{maxHK0},v_{maxmAb0},K_{mGLC_{0}},v_{maxg0}]=[4.15\times 10^{-5},3.5\times 10^{-8},4.2,0.126]$; (2) $\boldsymbol{\alpha}_{1}=[v_{maxHK1},v_{maxmAb1},K_{mGLC_{1}}]=[1.95\times 10^{-5},7.5\times 10^{-8},3.5]$ and (3) $\boldsymbol{\beta}=[\beta_{0},\beta_{1}]=[0.3,0.1]$.
@@ 6.2 Digital Twin Calibration
    With the true parameters defined in last subsection, our objective is to calibrate the digital twin model by strategically designing CHO cell culture experiments. The aim is to minimize the model prediction error by choosing the initial glucose concentration (design variable). In each experiment, we adjust the initial glucose concentration and then collect a new batch of data to continuously refine our model.
    For the calibration experiments setting, we assume time course data of the states is collected every 12h until 72h, so each batch includes 7-time sequence data. We focus on calibrating parameters $\boldsymbol{\alpha}_{0},\boldsymbol{\alpha}_{1}$. The parameter $\boldsymbol{\beta}$ is not considered during the digital twin calibration because it’s directly related to cell density, which is not built into our state. The design space for the initial concentrations of glucose is set within the range $[18,38]$ mM, and the initial concentration of lactate and mAb are $[LAC]_{0}=0.8$ mM, $[mAb]_{0}=0$ mM, the initial cell densities in Phase 0 and 1 are set to be $X_{0,0}=180\times 10^{6}$ cells/L, $X_{1,0}=10\times 10^{6}$ cells/L, respectively. The initial parameters for the digital twin are estimated by the first batch of data using MLE. Since our parameters are in different scales, we set a learning rate $\lambda=[3\times 10^{-5},3\times 10^{-7},3,0.3,3\times 10^{-5},3\times 10^{-7},0.3]$ for updating the parameter and $\gamma_{k}=5$ (for all $k$) for updating next design policy. The fixed set of test data $\mathcal{D}_{test}$ is composed of initial states $\boldsymbol{s}_{0}^{1}=[28,0.8,0]$ and $\boldsymbol{s}_{0}^{2}=[38,0.8,0]$.
    We conducted 15 macro-replications, each consisting of $K=200$ iterations under varying initial conditions, with initial glucose levels randomly chosen from the uniform distribution over the interval $[18,38]$. We compare the proposed gradient-based calibration approach with (1) a random design approach, where the design is uniformly sampled from $[8,38]$; (2) GP-based approach with expected improvement acquisition function and 20 initial points. A Matern Kernal $K_{\text{Matern}}\left(x,x^{\prime}\right)=\frac{2^{1-\nu}}{\Gamma(\nu)}\left(\frac{\sqrt{2\nu}|d|}{\ell}\right)^{\nu}K_{\nu}\left(\frac{\sqrt{2\nu}|d|}{\ell}\right)$ is used with $d=x-x^{\prime}$, $\ell=1$ and $\nu=1.5$. The GP regression is fitted using the MSE as the output and parameters as input. The GP model is iteratively updated with data from new physical experiments. To ensure consistency across the three approaches, we used the same random seeds for these three approaches. The mean prediction performance and the 95% confidence interval, calculated across macro-replications, represented by the Mean Relative Errors $\text{MRE}=\frac{1}{7}\sum_{j=1}^{7}{|\widehat{\theta}_{(j)}-\theta_{(j)}^{c}|}/{|\theta_{(j)}^{c}|}$ with script $(j)$ representing the $j$-th component, are presented in Figure 5 for mAb protein drug generation prediction and in Figure 5 for parameter estimation, respectively. For all figures, red lines represent the performance of our approach, green lines represent the random design and blue lines represent GP.
    Figure 4: mAb Prediction Error
@@
    The results in Figures 5 and 5 indicate that the proposed calibration approach is significantly more sample-efficient than the random design and GP-based calibration approach. The MRE for mAb production predictions using our method drops to about 60% after 60 iterations, compared to 220% with a random design. Furthermore, the MRE for our approach converges to about 35% after 60 iterations for parameter estimations, significantly lower than the 700% observed with the random policy. In contrast, the GP-based method requires around 125 iterations to reach similar convergence levels for mAb production predictions.
    Table 2: Comparison of runtime for various calibration methods
    Time (Hour) 50 Iterations 100 Iterations 200 Iterations 400 Iterations
    Proposed Approach 0.131 0.345 0.791 1.874
    Random Search 0.097 0.186 0.373 0.86
    Gaussian Process $0.195$ $0.866$ $5.695$ $>24$
    To assess the computational efficiency, the total runtimes are presented in Table 2. it indicates that the GP method’s computational time increases significantly with larger datasets due to its cubic computational complexity from matrix inversion (Williams and Rasmussen, 2006). In contrast, the proposed gradient method shows a sub-cubic and more manageable increase in computation time, indicating better scalability and efficiency for larger experiments.
@@ 7 CONCLUSION
    This study develops a robust calibration approach for a Bio-SoS mechanistic model in the context of cell culture processes. By strategically guiding more informative data collection through the proposed gradient-based calibration approach, we significantly enhance the fidelity and predictive accuracy of the digital twin model. The empirical validation using the CHO cell culture model underscores the superiority of our approach over the traditional random design, showcasing its potential applicability across various biological systems. Overall, this work not only contributes to the theoretical advancements in bioprocess digital twin development but also holds promise for practical implementations that could improve the efficiency and effectiveness of biomanufacturing in the pharmaceutical industry.